OpenStudy (anonymous):

Find the equation of a line given the point and slope below. Arrange your answer in the form y = mx + b, where b is the constant. (-1, 0) m = -2 i got y=-2

OpenStudy (anonymous):

y-0=-2(x+1) y=-2x-2 so b=-2 not y=-2

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

:|

OpenStudy (anonymous):

take the "b" you have solved for in every single one of these, and plug it into "y = mx + b" along with the slope they gave you. You are NOT trying.